OpenStudy (compassionate):

\frac{1}{x] = \frac{1}{6} = \frac{1}{4}

OpenStudy (compassionate):

\frac{1}{x] = \frac{1}{6} = \frac{1}{4}

OpenStudy (compassionate):

1/x = 1/6 + 1/4

OpenStudy (astrophysics):

ahah

OpenStudy (compassionate):

I tried the rule and failed 1 * 4 + 1 * 6/6* 4

OpenStudy (compassionate):

10/24 doesn't make much sense

OpenStudy (compassionate):

Unless we can simplify it or something

OpenStudy (astrophysics):

\[\frac{ 1 }{ x } = \frac{ 1 }{ 6 }+\frac{ 1 }{ 4 }\] this?

OpenStudy (compassionate):

\[\frac{ 1 * 6 + 1 * 4 }{ 6 * 4 } = \frac{ 10 }{ 24 }\]

OpenStudy (compassionate):

However, my book says \[\frac{ 2 + 3 }{ 12 }\]

OpenStudy (astrophysics):

remember you have 1/x =... not x = ...

OpenStudy (compassionate):

So.. I'm just finding the LCD of 10 and 24... Ah.

OpenStudy (astrophysics):

\[\frac{ 1 }{ x } = \frac{ 1 }{ 6 } + \frac{ 1 }{ 4 } \implies \frac{ 1 }{ x } = \frac{ 10 }{ 24 } \implies \frac{ 24 }{ 10 } = x\]

OpenStudy (astrophysics):

There we go, the answer should be x=12/5

OpenStudy (astrophysics):

which is same as 24/10

OpenStudy (compassionate):

\\[2\frac{ 2 }{ 5}\][\frac{ 12 }{ 5}\] Since 5 goes into 12 two times with a remainder of 2, I can say

OpenStudy (astrophysics):

Sure thing

OpenStudy (zzr0ck3r):

nm i see